if any question like this 


x=.........,    y=............


then how to find relation b/w x and y


 

Grade:

2 Answers

anil kumar
11 years ago

Those questions are somewhat difficult to solve but its not impossible If x and y r given in the same parameters then one has to eliminate that parameter to get relation between x and y For instance : if x = cos (@) Y = sin ( @ ) Then v have to eliminate the @ In this case the relation is too easy cos^2(@)+sin^2(@)=1 I.e x^2+y^2=1 Thats the required relation This is just basic kind of problem Where others can be like x=t^2+2t+1 and y = t^2 -2t +1 This one requires examining Which gives x = (t+1)^2 and y = (t-1)^2 That gives root(x)-root(y)=2 Which is required answer The more models u practice the more skilled u become and u can find the ans giving method

shivang belwariar
11 years ago

just divide x&y and solve it

 

also u can put sin on bot side on both cases and solve it

 

plz aproove my answer